A box contains 5 red balls, 6 white balls and 9 black balls. Two balls are drawn at random. (The first ball is not replaced). Find the probability that both balls are of the same colour.


1
Expert's answer
2021-05-12T15:50:51-0400

Number of red balls =5

Number of White balls =6

Number of Black balls =9


Total Number of balls =20


Probability that both balls are of the same color-

"P=\dfrac{^5C_2}{^{20}C_2}+\dfrac{^6C_2}{^{20}C_2}+\dfrac{^9C_2}{^{20}C_2}"


"=\dfrac{20}{380}+\dfrac{30}{380}+\dfrac{72}{380}\\[9pt]=\dfrac{122}{380}=0.3210"